ABC, Co. uses 315 boxes of file folders per year. The price is $7.25 per box for an order size of 249 boxes or less. For orders of 250 to 999, the price is $7.00 per box. Carrying cost is 20% of the unit price, and ordering costs are $10 per order.

How many boxes should they order each time (round to the nearest integer)?

A.

157

B.

250

C.

66

D.

57

A product has demand during lead time of 80 units, with a standard deviation of 20 units. What safety stock provides (approximately) a 95% service level?

A.

80

B.

20

C.

33

D.

44

Expected completion time for a project is 57 weeks and standard deviation is 4.2 weeks. Assume normal distribution. What is the probability that the project will be completed within 60 weeks of starting it?

A.

0.155

B.

0.762

C.

0.417

D.

0.603

Demand is normally distributed with an average of 42 tulips per day, with a standard deviation of 5 units. With a marginal loss of $1.85 and a marginal profit of $3.15, what should their daily stocking policy be? Round off your answer to the nearest integer.

A.

42

B.

43

C.

44

D.

45
Answer #1

Demand (D) = 315

Ordering cost (S) = 10

Holding cost (H) = Carrying cost %* Price = 20%*Price

Assuming price as 7.25

EOQ = sqrt(2*D*S/H) = sqrt(2*315*10/(7.25*20%)) = 66

Total cost = D/EOQ*S + EOQ/2*H + Price*D = 315/66*10 + 66/2*20%*7.25 + 7.25*315 = 2379

For Price = 7 and EOQ = 250

Total cost = 315/250*10 + 250/2*20%*7 + 7*315 = 2393

So, lowest cost is in 1st scenario of EOQ = 66

Ans C